OpenStudy (anonymous):

solve for y 3x+2y+40=x-y

OpenStudy (anonymous):

Well, when we solve for y, we want to isolate y by itself, and its coefficient should be positive 1. So, I will subtract 3x +40 from each side So we get: 3x +2y + 40 = x -y 3x+2y +40 -3x -40 = x - y - 3x - 40 2y = -2x -y -40 Adding y on each side: 2y + y = -2x -y -40 +y 3y = -2x -40 isolating y, by divinding by 3: y = (-2x -40) /3

OpenStudy (anonymous):

thanks you!